# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v2

@@ 4.3 Full Bayesian implementation using MCMC
For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters. Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 4.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform(0,1)(0,1) priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components ishwaran2001gibbs . Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP. The local distribution indicators $z_{ij}^{l}$, local distribution proportion $\pi^{l}$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates. However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ univariate stepping-out slice sampling neal2003slice steps. Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3. For signal detection, we use an FDR-controlled, FNR-optimized hypothesis testing framework based on posterior draws for $\lambda_{ij}$, described next.
@@ 5 SIMULATION ASSESSMENTS
In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings. We follow the SRS contingency table generation approach previously used in the literature chakraborty2022use ; liu2024pattern . Specifically, to generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns. Using a reference $I\times J$ contingency table (described next) from a real-world SRS database, we extract row and column totals $\{\bm{n}_{i\bullet}:i=1,\dots,I\}$ and $\{\bm{n}_{\bullet j}:j=1,\dots,J\}$, respectively, to facilitate random contingency table generation. Next, we generate $\bm{p}_{i\bullet}\sim\text{Dirichlet}(\bm{n}_{i\bullet})$ and $\bm{p}_{\bullet j}\sim\text{Dirichlet}(\bm{n}_{\bullet j})$ for all $i$, $j$ defining the random probability matrix $\bm{p}=((p_{ij}))$ with $p_{ij}\propto\lambda_{ij}^{0}\ \bm{p}_{i\bullet}\ \bm{p}_{\bullet j}$ and $\sum_{i,j}p_{ij}=1$. Finally, we generate a random SRS contingency table by sampling from $\text{Multinomial}(\bm{n}_{\bullet\bullet};\bm{p})$, collapsing any generated $n_{ij}=1$ corresponding to a true signal (as defined by $\lambda^{0}_{ij}$) to $n_{ij}=0$ afterward to eliminate signal detection at $n_{ij}=1$.
@@
We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,61063{,}976{,}610 reported cases chakraborty2023likelihood , to construct our reference SRS contingency table. We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category. For expository purposes, and to create a sufficiently large SRS contingency table, we focused on 1491 preferred terms (PTs) as AEs that are common (appear at least once) for these statin drugs as well for a separate group of drugs unrelated to statin, namely, Gadolinium-based contrast agents (GBCA) drugs chakraborty2022use . All other PTs recorded in the database within this period were collapsed into a heterogeneous category labeled “Other AE”, yielding a $1492\times 7$ contingency table summarizing the report counts for each drug-AE pair.
@@
To generate $\Lambda^{0}$, we define two signal patterns: (1) fixed row signals, where all drugs (columns) exhibit signals for the same AEs (PTs; rows), and (2) random signals, where signals appear on randomly chosen AEs (different row for each drug). A larger number of fixed row signals increase the between-drug associations, measured via correlations between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ for different statin drugs $j$ and $j^{\prime}$, while more random signals weaken them. The matrix $\Lambda^{0}$ is constructed by specifying fixed row signals and embedding additional random signals in the remaining rows across all statin columns. Signal strengths in these cells range from $1<\lambda_{ij}^{0}\leq 3$ (see Table 1). The remaining cells $(i^{\prime},j^{\prime})$ are designated as non-signals with $\lambda_{i^{\prime}j^{\prime}}^{0}=1$. To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (0 or $0.30$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$. This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables. Within this setup, we designed three simulation scenarios, each with and without zero-inflation (see Table 1).
@@
Simulation Scenario 1a 1b 2a 2b 3a 3b
No. of fixed row signals 5 5 10 10 20 20
No. of random signals per column 17 17 10 10 3 3
Zero-inflation rate per column 0 0.3 0 0.3 0 0.3
Average Kendall’s $\tau$ between columns 0.12 0.12 0.49 0.49 0.77 0.77
  • A common signal strength value $\lambda^{0,\text{signal}}$ is assigned to all signal cells in a given simulation setting, and we vary $\lambda^{0,\text{signal}}\in\{1.1,1.2,1.3,1.4,1.5,1.7,1.9,2,2.5,3\}$.
@@
We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al.hu2015signal (henceforth DP Hu et al). For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 4.4), with the additional constraint $n_{ij}>1$ for our methods. To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted muller2006fdr posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$. Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT chakraborty2022use ; huang2011likelihood ; ding2020evaluation , which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present.
@@
The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is very small ($\tau=0.12$; Cases 1a,b). Under moderate between-drug association ($\tau=0.49$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$). The gain is even further pronounced under high between-drug association ($\tau=0.77$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$. These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models. In all settings, the FDR and average type I error is strictly controlled for all methods.
@@
All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz. The median computation time per replication for our DP local-global methods (both Poisson and ZIP) is 48 minutes.
@@ 6 REAL DATA ANALYSIS
To further examine the signal detection results, we focus on our DP Poisson local-global model and zoom in on 15 selected PTs, whose signal strength estimates are plotted using heatmaps in Figure 4. The figure is divided into three panels based on the frequencies of the AEs, and the colors in the heatmaps indicate either non-detection (gray) or the estimated signal strengths measured through the posterior medians of $\lambda_{ij}$ (white-to-blue gradient with a darker blue signifying a stronger signal strength). Panel A displays the five most common AEs with the largest frequencies (total counts for the AEs across all six statin drugs being greater than $2000$) and the strongest estimated signal strengths (having the largest posterior medians of $\lambda_{ij}$ across all statin drugs $j$) that are detected as a signal for at least one statin drug by our method. The AEs displayed in this panel are all well-documented muscle-related conditions, such as rhabdomyolysis antons2006clinical ; tournadre2020statins and myopathy ballantyne2003risk ; abd2011statin , and their detection by our method is consistent with the performance of all competing methods. Panel B contains five semi-rare AEs (with observed counts across statin drugs ranging between 500 and 2000) that exhibit the strongest estimated signal strengths. This category primarily includes some specific types of disorders, such as nervous system disorders, and interestingly, several of these detected signals are not identified by any competing method for the newer or less frequently used statin drugs. For example, unlike other methods, our model detects lovastatin-induced drug intolerance as a signal, which has been flagged as a potential side effect of lovastatinahmad2014statin . Additionally, our method detects signals for paraesthesia and hypoaesthesia, symptoms of peripheral neuropathy, associated with lovastatin use; this finding is also consistent with previous research chong2004statin ; wang2021association . By appropriately borrowing strengths from the more commonly used statins, our method is successfully able to identify potential lovastatin-induced AE signals in these specific examples. Finally, Panel C focuses on five rare AEs (with observed counts across statin drugs being lower than 500) with the strongest estimated signal strengths. Within this group, our method identifies several AEs related to neurological and movement issues, such as movement disorder and swollen tongue, and some other mixed disorders. Many of these AEs were not originally discovered in clinical trials, but later were flagged as potential signals in specific, targeted pharmacovigilance investigations. For example, fluvastatin-induced swollen tongue, also known as angioedema, has been recently recorded kumar2024fluvastatin . Similar to the semi-rare AEs displayed in Panel B, our method is able to identify these rare AE signals for the less frequently used drug fluvastatin by appropriately sharing AE signal strength information across all statin drugs.

### v3

@@ 4.3 Full Bayesian implementation using MCMC
For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters. Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 4.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform(0,1)(0,1) priors for the probability weight parameters, viz., $\pi$ and $\omega_{j}$ (if included in the model).
We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components Ishwaran and James (2001). Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP. The local distribution indicators $z_{ij}$, local distribution proportion $\pi$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates. However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ univariate stepping-out slice sampling Neal (2003) steps. Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3. For signal detection, we use an FDR-controlled, FNR-optimized hypothesis testing framework based on posterior draws for $\lambda_{ij}$, described next.
@@ 5 SIMULATION ASSESSMENTS
In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings. We follow the SRS contingency table generation approach previously used in the literature Chakraborty et al. (2022); Liu et al. (2024). Specifically, to generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns. Using a reference $I\times J$ contingency table (described next) from a real-world SRS database, we extract row and column totals $\{\bm{n}_{i\bullet}:i=1,\dots,I\}$ and $\{\bm{n}_{\bullet j}:j=1,\dots,J\}$, respectively, to facilitate random contingency table generation. Next, we generate $\bm{p}_{i\bullet}\sim\text{Dirichlet}(\bm{n}_{i\bullet})$ and $\bm{p}_{\bullet j}\sim\text{Dirichlet}(\bm{n}_{\bullet j})$ for all $i$, $j$ defining the random probability matrix $\bm{p}=((p_{ij}))$ with $p_{ij}\propto\lambda_{ij}^{0}\ \bm{p}_{i\bullet}\ \bm{p}_{\bullet j}$ and $\sum_{i,j}p_{ij}=1$. Finally, we generate a random SRS contingency table by sampling from $\text{Multinomial}(\bm{n}_{\bullet\bullet};\bm{p})$, collapsing any generated $n_{ij}=1$ corresponding to a true signal (as defined by $\lambda^{0}_{ij}$) to $n_{ij}=0$ afterward to eliminate signal detection at $n_{ij}=1$.
@@
We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,61063{,}976{,}610 reported cases Chakraborty et al. (2023), to construct our reference SRS contingency table. We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category. For expository purposes, and to create a sufficiently large SRS contingency table, we focused on $I=1491$ preferred terms (PTs) as AEs that are common (appear at least once) for these statin drugs as well for a separate group of drugs unrelated to statin, namely, Gadolinium-based contrast agents (GBCA) drugs Chakraborty et al. (2022). All other PTs recorded in the database within this period were collapsed into a heterogeneous category labeled “Other AE”, yielding a $1492\times 7$ contingency table summarizing the report counts for each drug-AE pair. Alongside this larger setting, we also consider a smaller setting involving a subset of this dataset with $I=46$ relevant AEs as considered in prior workChakraborty et al. (2022); Huang et al. (2011) and a collapsed Other AE cateogory; see Supplement S.2.5.
@@
To generate $\Lambda^{0}$, we define two signal patterns: (1) fixed row signals, where all drugs (columns) exhibit signals for the same AEs (PTs; rows), and (2) random signals, where signals appear on randomly chosen AEs (different row for each drug). A larger number of fixed row signals increase the between-drug associations, measured via correlations between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ for different statin drugs $j$ and $j^{\prime}$, while more random signals weaken them. The matrix $\Lambda^{0}$ is constructed by specifying fixed row signals and embedding additional random signals in the remaining rows across all statin columns. Signal strengths in these cells range from $1<\lambda_{ij}^{0}\leq 3$ (see Table 1). The remaining cells $(i^{\prime},j^{\prime})$ are designated as non-signals with $\lambda_{i^{\prime}j^{\prime}}^{0}=1$. To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (0 or $0.30$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$. This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables. Within this setup, we consider four simulation scenarios with varied range of between-column correlations in $\Lambda^{0}$, each with and without zero-inflation (see Table 1).
@@
Simulation Scenario 0a 0b 1a 1b 2a 2b 3a 3b
No. of fixed row signals 1 1 3 3 10 10 20 20
No. of random signals per column 30 30 20 20 10 10 3 3
Zero-inflation rate per column 0 0.3 0 0.3 0 0.3 0 0.3
Average Kendall’s $\tau$ between column pairs $\leq$0.01 $\leq$0.01 0.12 0.12 0.49 0.49 0.77 0.77
  • A common signal strength value $\lambda^{0,\text{signal}}$ is assigned to all signal cells in a given simulation setting, and we vary $\lambda^{0,\text{signal}}\in\{1.01,1.1,1.2,1.3,1.4,1.5,1.7,1.9,2,2.5,3\}$.
@@
We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al.Hu et al. (2015) (henceforth DP Hu et al). For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 4.4), with the additional constraint $n_{ij}>1$ for our methods. To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted Müller et al. (2006) posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$. Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT Chakraborty et al. (2022); Huang et al. (2011); Ding et al. (2020), which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present.
@@
Our results show that across all settings, the FDR and average type I error are strictly controlled by all methods except BCPNN, which appears to entirely lose control of FDR under the near-null setting of $\lambda^{0,\text{signal}}=1.01$, despite the use of a Benjamini-Hochberg-based FDR adjustment during signal detection.
Focusing on sensitivity, the first four rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the there is very weak ($\tau\leq 0.01$; Cases 0a, b) or weak ($\tau=0.12$; Cases 1a,b) between-drug associations. Under moderate between-drug association ($\tau=0.49$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$). The gain is even further pronounced under high between-drug association ($\tau=0.77$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$. These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models.
@@
The results from the smaller setting involving $I=46$ AEs are essentially similar to those obtained in the larger setting; see Supplementary Figure S.2 in Supplement S.3.4 for further details. All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz. The median computation time per replication for our DP local-global methods (both Poisson and ZIP) is approximately 48 minutes for the larger setting with 1491 AEs, compared with 0.66, 0.02, 0.42, and 26 minutes for BCPNN, GPS, LRT, and DP Hu et al., respectively. For the smaller 46-AE dataset, our methods require about 2 minutes per replication, versus 0.02, 0.002, 0.07, and 0.7 minutes for the same competing approaches. The higher computational cost reflects our substantially richer modeling assumptions: except for DP Hu et al., the competing methods achieve highly efficient computation through partial analytical tractability but do not provide flexible, regularized estimation of signal strength parameters with full uncertainty quantification, nor do they capture complex between-drug associations in AE signal patterns. Relative to DP Hu et al., our methods introduce only one additional global DP component and the local-global mixture parameters, incurring modest extra cost while delivering substantial gains in sensitivity and signal detection, as demonstrated in our simulations.
Remark.
Leveraging the simulation setup described as above, we performed additional sensitivity analyses to examine the robustness of the our methods with respect to the choice of hyperparameters, including the prior hyperparameters $\psi_{\alpha}$ and $\psi_{\beta}$, as well as the grid specification for the hypothesis testing threshold $k$. For each of $\psi_{\alpha}$ and $\psi_{\beta}$, we considered two alternative choices representing values toward the lower and upper extremes of the weakly informative range along with the default $\psi_{\alpha}=3$ and $\psi_{\beta}=1/2$. Specifically, for $\psi_{\alpha}$ we considered $\psi_{\alpha}=1$ (placing $\approx 64\%$ mass in $(0,1)$) and $\psi_{\alpha}=5$ ($\approx 99\%$ mass in $(0,1)$), and for $\psi_{\beta}$ we considered $\psi_{\beta}=1/7$ ($\approx 90\%$ mass in $(0,1)$) and $\psi_{\beta}=6$ (only $10\%$ mass in $(0,1)$). For the $k$-grid construction, we considered a more adaptive approach monitoring the computed maximum of all $\lambda_{ij}$ posterior draws instead of the default grid on $[1.1,3]$. Across all settings, the resulting performance curves were virtually identical to those obtained under the default specifications, indicating that the proposed method is largely stable across a range of reasonable prior choices and grid settings. A detailed description of the sensitivity analysis results is provided in the Supplement S.3.2.
@@ 6 REAL DATA ANALYSIS
To further examine the signal detection results, we focus on our DP Poisson local-global model and zoom in on 15 selected PTs, whose signal strength estimates are plotted using heatmaps in Figure 4. The figure is divided into three panels based on the frequencies of the AEs, and the colors in the heatmaps indicate either non-detection (gray) or the estimated signal strengths measured through the posterior medians of $\lambda_{ij}$ (white-to-blue gradient with a darker blue signifying a stronger signal strength). Panel A displays the five most common AEs with the largest frequencies (total counts for the AEs across all six statin drugs being greater than $2000$) and the strongest estimated signal strengths (having the largest posterior medians of $\lambda_{ij}$ across all statin drugs $j$) that are detected as a signal for at least one statin drug by our method. The AEs displayed in this panel are all well-documented muscle-related conditions, such as rhabdomyolysis Antons et al. (2006); Tournadre (2020) and myopathy Ballantyne et al. (2003); Abd and Jacobson (2011), and their detection by our method is consistent with the performance of all competing methods. Panel B contains five semi-rare AEs (with observed counts across statin drugs ranging between 500 and 2000) that exhibit the strongest estimated signal strengths. This category primarily includes some specific types of disorders, such as nervous system disorders, and interestingly, several of these detected signals are not identified by any competing method for the newer or less frequently used statin drugs. For example, unlike other methods, our model detects lovastatin-induced drug intolerance as a signal, which has been flagged as a potential side effect of lovastatinAhmad (2014). Additionally, our method detects signals for paraesthesia and hypoaesthesia, symptoms of peripheral neuropathy, associated with lovastatin use; this finding is also consistent with previous research Chong et al. (2004); Wang et al. (2021). By appropriately borrowing strengths from the more commonly used statins, our method is successfully able to identify potential lovastatin-induced AE signals in these specific examples. Finally, Panel C focuses on five rare AEs (with observed counts across statin drugs being lower than 500) with the strongest estimated signal strengths. Within this group, our method identifies several AEs related to neurological and movement issues, such as movement disorder and swollen tongue, and some other mixed disorders. Many of these AEs were not originally discovered in clinical trials, but later were flagged as potential signals in specific, targeted pharmacovigilance investigations. For example, fluvastatin-induced swollen tongue, also known as angioedema, has been recently recorded Kumar et al. (2024). Similar to the semi-rare AEs displayed in Panel B, our method is able to identify these rare AE signals for the less frequently used drug fluvastatin by appropriately sharing AE signal strength information across all statin drugs.